OpenStudy (babyslapmafro):

I am asked to find the vector component of v along b and the vector orthogonal to b. Could someone please help provide formulas to solve these problems or just help in general? Thanks...

OpenStudy (anonymous):

Is that a linear algebra problem? Have you learned the inner product?

OpenStudy (babyslapmafro):

Calc III

OpenStudy (anonymous):

Say the vector that is orthogonal to b is x. Then the inner/dot product of b and x will be zero. I don't know if this can help.

OpenStudy (loser66):

@giolefpatceid it applies linear algebra to cal 3, the same formula

OpenStudy (loser66):

\[proj_a u=\frac{u(dot)a}{||a||^2}a\]

OpenStudy (babyslapmafro):

which component does that formula provide?

OpenStudy (babyslapmafro):

It provides the orthogonal projection, yes? How do I find the vector component of v along b then?

OpenStudy (babyslapmafro):

@loser66

OpenStudy (loser66):

@e.mccormick explain him, please

OpenStudy (loser66):

from my formula, it calculate vector u along a , you just replace u =v and a =b so you have vector v along b, does it make sense?

OpenStudy (babyslapmafro):

ok then how do I find the vector component orthogonal to b?

OpenStudy (e.mccormick):

A projection is short for Orthogonal Projection.

OpenStudy (e.mccormick):

This has a good example of how that formula is made and a diagram of what it represents: http://www.math4all.in/public_html/linear%20algebra/chapter8.1.html

OpenStudy (babyslapmafro):

@loser my book says that the formula you provided me defines the orthogonal projection of v on vector b.

OpenStudy (e.mccormick):

Does your book talk about subtracting the projection from the vector to get the component that is orthogonal?

OpenStudy (babyslapmafro):

ok yes I got it now...the way my book explained this is a bit misleading to me...thanks for the help.

OpenStudy (e.mccormick):

Something like this? Find the vector component \(\vec{w}\) of \(\vec{u}\) orthogonal to \(\vec{v}\): \( w = u - \cfrac{{u \cdot v}}{{v \cdot v}}v\)

OpenStudy (e.mccormick):

If you ever run into the Ghram-Schmidt Orthornormalization, you will see some of this pop up again.

OpenStudy (babyslapmafro):

ok thanks